(Fatoring Trinomials of the type ax squared + bx + c)
Factor each experession.
1. 5x squared - 2x - 7
2. 7n squared + 9n + 2
3. 3c squared - 17c - 6
4. 3x squared + 8x + 4

Answer by checkley71(8403)

5X^2-2X-7=(5X-7)(X+1)
7N+9N+2=(7N+2)(N+1)
3C^2-17C-6=(3C+1)(C-6)
3X^2+8X+4=(3X+2)(X+2)